Suppose your firm is considering investing in a project with the cash flows shown below, that the required rate of return on projects of this risk class is 11%, and that the maximum allowable payback and discounted payback statistics for your company are 3 and 3.5 years, respectively.

Time 0 1 2 3 4 5
Cash flow -$235,000 $65,800 $84,000 $141,000 $122,000 $81,200

Use the NPV decision rule to evaluate this project; should it be accepted or rejected?

Use Excel and show work when completing.
Answer #1

A949 А 926 Time Cash flow Cumulative cash flow Present value calculation Present value calculation Cumulative present value c

Time Cash flow Cumulative cash flow Present value calculation Present value calculation Cumulative present value cash flow
0 -2,35,000.00 -2,35,000.00 -235000 -2,35,000.00 -2,35,000.00
1 65,800.00 -1,69,200.00 =65800/(1+11%)^1 59,279.28 -1,75,720.72
2 84,000.00 -85,200.00 =84000/(1+11%)^2 68,176.28 -1,07,544.44
3 1,41,000.00 55,800.00 =141000/(1+11%)^3 1,03,097.98 -4,446.45
4 1,22,000.00 1,77,800.00 =122000/(1+11%)^4 80,365.18 75,918.73
5 81,200.00 2,59,000.00 =81200/(1+11%)^5 48,188.25 1,24,106.98
IRR Excel formula =IRR(B927:B932)
IRR 28.79%
NPV Excel formula =NPV(11%,B928:B932)+B927
NPV 1,24,106.98
Payback period calculation =2+(85200/141000)
Payback period 2.60
Discounted payback period calculation =3+(4446.45/80365.18)
Payback period 3.06
This project should be accepted as it's IRR is higher than required rate, NPV is positive and 124,106.98, Its payback and discounted payback periods are 2.60 years and 3.06 years respectively which are below the maximum allowable payback and discounted payback period.